OpenStudy (anonymous):

If a cylinder’s radius and height are each shrunk down to a third of the original size, what would be the formula to find the modified surface area?

OpenStudy (anonymous):

A.\[\frac{ 1 }{ 9 } \pi r^2 + \frac{ 2 }{ 3 } \pi r h\] B.\[\frac{ 2 }{ 3 } \pi r^2 + \frac{ 2 }{ 3 } \pi r h\] C.\[\frac{ 2 }{ 3 } \pi r^2 + \frac{ 1 }{ 3 } \pi r h\] D.\[\frac{ 2 }{ 9 } \pi r^2 + \frac{ 2 }{ 9 } \pi r h\]

OpenStudy (mathstudent55):

What is the formula for the surface area of a cylinder?

OpenStudy (anonymous):

\[2 \pi r h + 2 \pi r^2\]

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Correct. Now use the formula you have and replace r with r/3, and replace h with h/3. Then simplify.

OpenStudy (anonymous):

so itd be B?

OpenStudy (mathstudent55):

I don't know until I try. Let's see.

OpenStudy (mathstudent55):

Original formula: \(SA = 2 \pi r h + 2 \pi r^2\) No let r = r/3, and h = h/3: \(SA = 2 \pi \dfrac{r}{3} h + 2 \pi \left(\dfrac{r}{3} \right) ^2\) \(SA = \dfrac{2}{3} \pi r h + \dfrac{2}{9} \pi r^2\)

OpenStudy (mathstudent55):

Did you copy all your choices correctly?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

none of the choices are correct.

OpenStudy (anonymous):

it was d

OpenStudy (mathstudent55):

Then you copied d incorrectly above. Only one fraction has a denominator of 9, not both. Since the radius is squared in the formula of the area of a circle, the part of the surface area that comes from the bases has a denominator of 9 since (r/3)^2 = r^2/9. The lateral area uses the circumference of the base. When the radius becomes 1/3 of the radius, you just use r/3 in the circumference formula. There is no squaring of the radius in the circumference formula, so the denominator remains 3, not 9.